# Balanced Accuracy: The Right Metric for Evaluating LLM Judges Explained through Youden’s J statistic

### v1

@@ Abstract
Explained through Youden’s J statistic
@@ 1 Introduction
Evaluating large language models (LLMs) is a cornerstone of their development cycle. Standard practice involves running models on benchmark datasets of user prompts and estimating the prevalence of key behaviors in their responses such as task pass rates, safety violations, or false refusals. Prevalence estimates rely on another classifier, typically an LLM, a fine-tuned model, or human annotators. We refer to this classifier as a judge li2024from; zheng2023judging. Because prevalence measurements feed directly into ablation studies, capabilities assessments, and release decisions, the quality of this judge critically determines the validity of the resulting model comparisons.
@@
We argue instead for Balanced Accuracy (equivalently, Macro-Recall) as the primary metric for judge selection. Balanced Accuracy is independent of class prevalence, assigns equal importance to both classes, extends naturally to multi-class settings, and most directly captures the key property needed for prevalence comparison: how well a judge distinguishes positive from negative instances. We formalize this by grounding the argument in Youden’s $J$ statistic youden1950index, historically used in diagnostic testing to measure a classifier’s ability to separate classes. We show that $J$ is theoretically aligned with detecting prevalence differences and that Balanced Accuracy is a simple monotonic linear transformation of it. We provide geometric intuition through ROC analysis and empirical examples demonstrating that Balanced Accuracy leads to more reliable judge selection and more trustworthy downstream evaluation.
@@ 2 Preliminaries
This work focuses on pointwise evaluation, where each model response is judged independently. This differs from pairwise evaluation, where responses are compared head-to-head; pairwise settings produce inherently balanced labels, making metrics like Accuracy suitable for evaluating preference models malik2025rewardbench2advancingreward.
@@ 3 The Pitfalls of Traditional Metrics
The same reasoning applies to AUC metrics: PR-AUC inherits Precision’s prevalence dependence, while ROC-AUC — built from Sensitivity and Specificity — is prevalence-independent and label-symmetric, making it more relevant for our task.
@@ 3.4 Issues with Accuracy
Accuracy, defined as $\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$, is perhaps the most widely used metric. In heavily imbalanced golden sets, which are common in safety evaluations, for example, a classifier can achieve high Accuracy simply by predicting the majority class. This makes Accuracy ill-suited for our goal of selecting judges who perform well across both classes.
@@ 3.5 Issues with Agreement Metrics
Agreement measures such as Cohen’s kappa cohen1960coefficient, Scott’s Pi scott1955reliability, and Krippendorff’s alpha krippendorff2004content are frequently used in human annotation settings, but they measure the wrong quantity for our purposes.
@@ 4 Youden’s J: A Metric Theoretically Aligned
The shortcomings of traditional metrics suggest that we should instead evaluate judges using a measure that (i) is independent of class prevalence, (ii) treats both classes symmetrically, and (iii) reflects how well the judge preserves true differences in prevalence between models. Youden’s $J$ statistic powers2011evaluation satisfies all three criteria and emerges naturally from the structure of the prevalence-estimation problem. Youden’s $J$ is defined for any binary classifier as:
@@ 4.2 Relationship to ROC Analysis
Youden’s $J$ also has a clear geometric interpretation in ROC space fawcett2006introduction. Each threshold of a classifier corresponds to a point (FPR, TPR) on the ROC curve, and $J$ is the vertical distance between this point and the diagonal chance line ($y=x$). (Figure 2).
@@
Remarks on calibration: When calibrating the decision threshold, we should select the threshold that maximizes Youden’s $J$. This requires labeled data to tune the threshold in addition to the golden set to prevent overfitting. In this paper we estimate prevalence by taking the mean of binary labels. Alternatively, one could estimate prevalence by taking the mean of the predicted score (a less common practice), in which case a metric like the Kuiper statistic kuiper1960tests, which assesses calibration across the entire range of predicted probabilities, is relevant.
In empirical analyses on our datasets, we observe that $J$ is highly correlated with ROC AUC (0.88), consistent with their shared reliance on TPR and FPR. By contrast, the correlation between $J$ and F1 is substantially weaker (0.56), reflecting F1’s dependence on Precision and hence on class prevalence. These observations support the theoretical distinction between prevalence-independent and prevalence-dependent metrics.
@@ 7 Conclusion
First, Balanced Accuracy summarizes performance through TPR and TNR, but it does not fully characterize a judge’s error profile. A judge with a higher Balanced Accuracy may still have lower precision, higher false-positive rates, or other operational characteristics that matter for specific applications (e.g., safety teams may prioritize very high recall on violation classes). As with any scalar metric, Balanced Accuracy should be complemented with inspection of the underlying confusion matrix and class-specific error rates.
Second, our analysis assumes that a judge’s error rates (TPR and FPR) remain stable across the assistant models being compared. If a judge exhibits model-specific bias such as the self-preference and family-preference effects observed when LLM judges evaluate outputs from related models Wataoka2024SelfPreference; zheng2023judging, then prevalence estimates may be distorted regardless of the metric used. Detecting and mitigating this form of judge bias is an essential but orthogonal challenge that must accompany any metric-based judge selection.

### v2

@@ Abstract
Explained through Youden’s J statisticthanks: A preprint of this work is available on arXiv.
@@ 1 Introduction
Evaluating large language models (LLMs) is a cornerstone of their development cycle. Standard practice involves running models on benchmark datasets of user prompts and estimating the prevalence of key behaviors in their responses such as task pass rates, safety violations, or false refusals. Prevalence estimates rely on another classifier, typically an LLM, a fine-tuned model, or human annotators. We refer to this classifier as a judge Gu et al. (2024); Liu et al. (2023); Li et al. (2024b, a); Zheng et al. (2023). Because prevalence measurements feed directly into ablation studies, capabilities assessments, and release decisions, the quality of this judge critically determines the validity of the resulting model comparisons.
@@
We argue instead for Balanced Accuracy (equivalently, Macro-Recall) as the primary metric for judge selection. Balanced Accuracy is independent of class prevalence, assigns equal importance to both classes, extends naturally to multi-class settings, and most directly captures the key property needed for prevalence comparison: how well a judge distinguishes positive from negative instances. We formalize this by grounding the argument in Youden’s $J$ statistic Youden (1950), historically used in diagnostic testing to measure a classifier’s ability to separate classes. We show that $J$ is theoretically aligned with detecting prevalence differences and that Balanced Accuracy is a simple monotonic linear transformation of it. We provide geometric intuition through ROC analysis and empirical examples demonstrating that Balanced Accuracy leads to more reliable judge selection and more trustworthy downstream evaluation.
@@ 2 Preliminaries
This work focuses on pointwise evaluation, where each model response is judged independently. This differs from pairwise evaluation, where responses are compared head-to-head; pairwise settings produce inherently balanced labels, making metrics like Accuracy suitable for evaluating preference models Malik et al. (2025).
@@ 3 The Pitfalls of Traditional Metrics
The same reasoning applies to AUC metrics: PR-AUC inherits Precision’s prevalence dependence, while ROC-AUC — built from Sensitivity and Specificity — is prevalence-independent and label-symmetric, making it more relevant for our task. PR-based summaries can also exhibit high statistical uncertainty under small test sets, particularly in the low-recall regime where the joint uncertainty becomes highly non-linear Urlus et al. (2023).
@@ 3.4 Issues with Accuracy
Accuracy, defined as $\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$, is perhaps the most widely used metric. In heavily imbalanced golden sets, which are common in safety evaluations, for example, a classifier can achieve high Accuracy simply by predicting the majority class. This makes Accuracy ill-suited for our goal of selecting judges who perform well across both classes Dorner et al. (2025b).
@@ 3.5 Issues with Agreement Metrics
Agreement measures such as Cohen’s kappa Cohen (1960), Scott’s Pi Scott (1955), and Krippendorff’s alpha Krippendorff (2004) are frequently used in human annotation settings. Similarly, correlation-based metrics such as Pearson correlation are used to measure alignment between judge scores and human ratings in pointwise rubric-based evaluation Kim et al. (2024). However, these metrics measure the wrong quantity for our purposes.
@@ 4 Youden’s J: A Metric Theoretically Aligned
The shortcomings of traditional metrics suggest that we should instead evaluate judges using a measure that (i) is independent of class prevalence, (ii) treats both classes symmetrically, and (iii) reflects how well the judge preserves true differences in prevalence between models. Youden’s $J$ statistic Powers (2011) satisfies all three criteria and emerges naturally from the structure of the prevalence-estimation problem. Youden’s $J$ is defined for any binary classifier as:
@@ 4.2 Relationship to ROC Analysis
Youden’s $J$ also has a clear geometric interpretation in ROC space Fawcett (2006). Each threshold of a classifier corresponds to a point (FPR, TPR) on the ROC curve, and $J$ is the vertical distance between this point and the diagonal chance line ($y=x$). (Figure 2).
@@
Remarks on calibration: When calibrating the decision threshold, we should select the threshold that maximizes Youden’s $J$. This requires labeled data to tune the threshold in addition to the golden set to prevent overfitting. In this paper we estimate prevalence by taking the mean of binary labels. Alternatively, one could estimate prevalence by taking the mean of the predicted score (a less common practice), in which case a metric like the Kuiper statistic Kuiper (1960), which assesses calibration across the entire range of predicted probabilities, is relevant.
In empirical analyses on our datasets, we observe that $J$ is highly correlated with ROC AUC (0.88), consistent with their shared reliance on TPR and FPR. By contrast, the correlation between $J$ and F1 is substantially weaker (0.56), reflecting F1’s dependence on Precision and hence on class prevalence. These observations support the theoretical distinction between prevalence-independent and prevalence-dependent metrics. More broadly, recent work has shown that the geometry of the ROC curve also determines the scaling behavior of test-time methods such as Best-of-N and rejection sampling when using LLM judges as verifiers Dorner et al. (2025a).
@@ 7 Conclusion
First, Balanced Accuracy summarizes performance through TPR and TNR, but it does not fully characterize a judge’s error profile. A judge with a higher Balanced Accuracy may still have lower precision, higher false-positive rates, or other operational characteristics that matter for specific applications (e.g., safety teams may prioritize very high recall on violation classes), while statistical debiasing approaches require high correlation between judge and ground truth scores to provide meaningful sample efficiency gains dorner2024limits. As with any scalar metric, Balanced Accuracy should be complemented with inspection of the underlying confusion matrix and class-specific error rates.
Second, our analysis assumes that a judge’s error rates (TPR and FPR) remain stable across the assistant models being compared. If a judge exhibits model-specific bias such as the self-preference and family-preference effects observed when LLM judges evaluate outputs from related models Wataoka et al. (2024); Zheng et al. (2023); Dorner et al. (2025b); Chen et al. (2024), then prevalence estimates may be distorted regardless of the metric used. Detecting and mitigating this form of judge bias is an essential but orthogonal challenge that must accompany any metric-based judge selection.